The unit of length convenient on the atomic scale is known as an angstrom and is
denoted by Å: 1 Å = 10.10 m. The size of a hydrogen atom is about 0.5 Å. What is the
total atomic volume in m3 of a mole of hydrogen atoms
Grade:11

2 Answers

Kevin Nash
10 years ago

Radius of hydrogen atom, r = 0.5 = 0.5 × 10–10 m

Volume of hydrogen atom =

 

1 mole of hydrogen contains 6.023 × 1023 hydrogen atoms.

∴ Volume of 1 mole of hydrogen atoms = 6.023 × 1023 × 0.524 × 10–30

= 3.16 × 10–7 m3

saji
6 years ago
4/3×3.14×(0.5×10^-10)^3=5.234×10^-31m^3Number of 1mole of H=Avogadro`s number(N)=6.023×10^235.236×10^-31×6.023×10^23m^3=3.152×10^-7m